# Generalization Below the Edge of Stability: The Role of Data Geometry

### v1

@@ 3.2 A Spectrum of Generalization on Isotropic Distributions
To explore the transition from a generalizing to a memorizing regime, we now analyze a family of isotropic distributions parameterized by a term that controls the concentration of data mass near the boundary of the unit ball. This allows for a precise characterization of how generalization degrades as data points become more radially exposed.
@@
Let $\bm{X}$} be a $d$-dimensional random vector in $\mathbb{R}^{d}$. For any $\alpha\in(0,\infty)$, the isotropic $\alpha$-powered-radial distribution is defined by the generation process
${\bm{X}}=h(R){\bm{U}}\sim\mathcal{P}_{{\bm{X}}}(\alpha),$ (8)
Note that as $\alpha\rightarrow 0$, the distribution $\mathcal{P}(\alpha)$ will be closer to the uniform distribution on the sphere.
@@
This framework enables the derivation of a generalization upper bound that depends explicitly on the parameter $\alpha$. The proof of the following theorem can be found in Appendix G.
Theorem 3.5.
@@
and for where $M\coloneqq\max\{D,\|f_{\bm{\theta}}|_{\mathbb{B}_{1}^{d}}\|_{L^{\infty}},1\}$ and $\lessapprox_{d}$ hides constants (which could depend on $d$) and logarithmic factors in $n$ and $(1/\delta)$.
To demonstrate the tightness of this result, we establish a corresponding lower bound. For this purpose, we consider the class of neural networks with bounded $g$-weighted path norm ${\bm{\Theta}}_{g}(\mathbb{B}_{1}^{d};1,1)$, where $g$ is the population version of the weighted.
@@
Let $\mathcal{P}$ denote any joint distribution of $({\bm{x}},y)$ where the marginal distribution of $\bm{x}$} is $\mathcal{P}_{{\bm{X}}}(\alpha))$ and $y$ is supported on [1,1][-1,1]. Let $\mathcal{D}_{n}=\{({\bm{x}}_{j},y_{j})\}_{j=1}^{n}$ be a dataset of $n$ i.i.d. samples from $\mathcal{P}$. Let $\widehat{R}_{\mathcal{D}_{n}}(f)$ be any empirical risk estimator for the true risk $R_{\mathcal{P}}(f):=\mathbb{E}_{({\bm{x}},y)\sim\mathcal{P}}[(f({\bm{x}})-y)^{2}]$. Then,
@@
Crucially, the proof is constructive and, as detailed in Appendix H, leverages the “neural shattering” observation found by (Liang et al., 2025). The method involves constructing a large family of distinct two-layer ReLU networks, and then showing that with high probability, this family contains at least one pair of networks that are indistinguishable to any learning algorithm that only has access to the training data. This construction exhibit how data geometry (more precisely, shatterability) connect to statistical uncertainty: (1) For bounded isotropic distributions with fixed boundary concentration (fixed $\alpha$) in high dimensions, the input space offers an exponential number of distinct directions; (2) any training set of finite size $n$ can only cover a fraction of these directions, leaving vast regions of the input space unsampled. Our proof exploits this by designing networks whose neurons activate only in localized, disjoint regions near the boundary of the unit ball. The construction ensures that, with high probability, a significant number of these regions contain no training data. By having the two networks differ only in these empty regions, they become identical on every point of the training set, yet remain substantially different in their population risk.
3.3 Flat Interpolation of Spherically Supported Data
The previous analysis indicates that generalization degrades as data concentrates toward a boundary. We now investigate the limiting case of this phenomenon, where the data support is confined to the unit sphere. In this setting, we show that the stability condition at the EoS is insufficient to prevent the network from perfectly interpolating the training data.
@@ 3.3 Flat Interpolation of Spherically Supported Data
Assume that $\{({\bm{x}}_{i},y_{i})\}_{i=1}^{n}$ is a dataset with ${\bm{x}}_{i}\in\mathbb{S}^{d-1}$ and pairwise distinct inputs, there exists a width $K\leq n$ network of the form (1) that interpolates the dataset and whose Hessian operator norm satisfies
@@
Both Theorem 3.7 our work and (Wen et al., 2023) construct explicit memorization networks to illustrate a negative result: even under natural “flatness/stability” proxies, a ReLU network can interpolate while failing to generalize. In (Wen et al., 2023), they exhibit trace-minimizing (“flattest”) interpolators that perform poorly on the distribution $\mathcal{P}_{\mathrm{xor}}$, which is different from our setting.
The existence of an interpolating network that remains stable demonstrates that the EoS condition alone does not preclude memorization. This result establishes a clear boundary for the effectiveness of the stability-induced implicit bias, highlighting that its success is critically dependent on the properties of the data’s support.
@@ 4.1 Empirical Verification of the Generalization Upper Bounds
(a) Adaptation to intrinsic dimension (b) Radial concentration parameter $\alpha$
Figure 2: How data geometry controls generalization. (a) Union of $J=20$ lines ($m=1$) embedded in $\mathbb{R}^{d}$ with $d\in\{10,50,100,500\}$. The regression slopes remain nearly constant across different $d$, showing that generalization adapts to intrinsic rather than ambient dimension. (b) Fixed ambient dimension $d=5$ with isotropic Beta-radial distributions (Definition G.1) for $\alpha\in\{1,5,10,50\}$. Larger $\alpha$ yields steeper slopes in the log–log error curve, consistent with improved rates as probability mass concentrates away from the boundary.
@@ 4.2 How Data Geometry Affects Representation Learning
We study how data geometry shapes the representation selected by GD at the BEoS regime through data activation rate of neurons. Given a neuron $v_{k}\phi({\bm{w}}_{k}^{\mathsf{T}}{\bm{x}}-b_{k})$ in the neuron network, its data activation rate is define to $\frac{1}{n}\sum_{i=1}^{n}\mathds{1}\{{\bm{w}}_{k}^{\top}{\bm{x}}_{i}>b_{k}\,\}$, which is exactly the probability term in the definition of the weight function $g$ in (25). Low data activation rate means the neuron fires on a small portion of the data.
@@
To verify it empirically, we compare two input distribution in $\mathbb{R}^{50}$: (i) uniform distribution on a sphere and (ii) a union of 20 lines by training ReLU networks the same recipe and initialization . As a result, the ReLU network trained on the sphere interpolates the noisy label quickly with final true MSE 1.0249 $\approx$ noise level ($\sigma^{2}=1$), while the the ReLU network trained on a union of lines resist to overfitting with final true MSE $0.07\approx 0$ (more details appears in Appendix C). Notably, the trained representations are presented in Figure 3. In particular, GD empirically finds our lower bound construction below the edge of stability.

### v2

@@ 3.1 A Spectrum of Generalization on Isotropic Distributions
We now specialize to isotropic distributions with compact support and assume that the maximal support radius is $1$. In this case, the depth of a point depends only on its radius. More precisely, we consider distributions of the form ${\bm{X}}=h(R)\,{\bm{U}},\quad{\bm{U}}\sim\mathrm{Uniform}(\mathbb{S}^{d-1})$, where $h$ is a radial profile and $R$ is a scalar random variable. After rescaling, we may assume the support of $\mathcal{P}_{X}$ is contained in $\mathbb{B}_{1}^{d}$. Writing $r=\|{\bm{x}}\|_{2}$ and setting $\varepsilon:=1-r$, we decompose $\mathbb{B}_{1}^{d}$ into (1) $\varepsilon$-annulus $\mathbb{A}^{d}_{\varepsilon}\coloneqq\{{\bm{x}}\in\mathbb{B}_{1}^{d}\mid\|{\bm{x}}\|_{2}\geq 1-\varepsilon\}$, (2) $\varepsilon$-strict interior $\mathbb{I}^{d}_{\varepsilon}\!:=\mathbb{B}_{1-\varepsilon}^{d}=\overline{\mathbb{B}^{d}_{1}\setminus\mathbb{A}_{\varepsilon}^{d}}$.
[…]
$\sup_{{\bm{\theta}}\in{{\bm{\Theta}}_{\mathrm{BEoS}}}(\eta,\mathcal{D})}\mathrm{Gap}(f_{{\bm{\theta}}},\mathcal{D})\;\leq\;\tilde{O}\Big(\mathcal{P}_{X}(\mathbb{A}_{\varepsilon}^{d})\Big)\;+\;\tilde{O}\Big(g_{\mathcal{P}}(1-\varepsilon)^{-\frac{d}{2d+3}}\,n^{-\frac{d+3}{4d+6}}\Big),\quad\text{w.h.p.}$ (7)
We instantiate this decomposition on a flexible family of radial profiles that interpolate between a “thick” center-concentration ball and a thin spherical shell.
@@
Let $\bm{X}$} be a $d$-dimensional random vector in $\mathbb{R}^{d}$. For any $\alpha\in(0,\infty)$, the isotropic $\mathrm{Beta}(\alpha)$-radial distribution is defined by
${\bm{X}}=h(R){\bm{U}}\sim\mathcal{P}_{X}(\alpha),$ (8)
where $R\sim\mathrm{Uniform}[0,1]$, ${\bm{U}}\sim\mathrm{Uniform}(\mathbb{S}^{d-1})$, and $h(r)=1-(1-r)^{1/\alpha}$ is a radial profile.
@@
Fix $\alpha\in(0,\infty)$. Let $\mathcal{P}(\alpha)$ be a joint distribution over $\mathbb{R}^{d}\times\mathbb{R}$ whose marginal over $\bm{x}$} is $\mathcal{P}_{{\bm{X}}}(\alpha)$. The corresponding labels $y$ are generated from a conditional distribution $\mathcal{P}(y|{\bm{x}})$ and are bounded: $|y|\leq D$ for some constant $D>0$.
This framework enables a generalization upper bound that depends explicitly on the parameter $\alpha$.
Theorem 3.4 (Spectrum of generalization on isotropic Beta-radial distributions).
@@
$\mathrm{Gap}_{\mathcal{P}}(f_{{\bm{\theta}}};\mathcal{D})\;\lessapprox_{d}\;\begin{cases}\Big(\frac{1}{\eta}-\frac{1}{2}+4M\Big)^{\frac{\alpha d}{d^{2}+4d+3}}\,M^{\frac{2d^{2}+7\alpha d+6\alpha}{d^{2}+4\alpha d+3\alpha}}\,n^{-\frac{\alpha(d+3)}{2(d^{2}+4\alpha d+3\alpha)}},&\alpha\geq\frac{3d}{2d-3},\\[3.99994pt] \Big(\frac{1}{\eta}-\frac{1}{2}+4M\Big)^{\frac{\alpha d}{d^{2}+4d+3}}\,M^{\frac{2d^{2}+7\alpha d+6\alpha}{d^{2}+4\alpha d+3\alpha}}\,n^{-\frac{\alpha}{2d+4\alpha}},&\alpha<\frac{3d}{2d-3},\end{cases}$ (9)
where $M\coloneqq\max\{D,\|f_{\bm{\theta}}|_{\mathbb{B}_{1}^{d}}\|_{L^{\infty}},1\}$ and $\lessapprox_{d}$ hides constants (which may depend on $d$) and logarithmic factors in $n$ and $(1/\delta)$.
Sketch proof of Theorem 3.4. Suppose $\mathcal{P}_{X}$ is a $\mathrm{Beta}(\alpha)$-radial distribution $\mathcal{P}_{X}(\alpha)$ with radial profile $h(r)=1-(1-r)^{1/\alpha}$ (Definition 3.2). Then Lemma G.2 gives $\mathbb{P}_{{\bm{X}}}(\mathbb{A}_{\varepsilon}^{d})\asymp\varepsilon^{\alpha}$ and Proposition G.6 gives $g_{\mathcal{P}}(1-\varepsilon)\asymp\varepsilon^{d+2\alpha}$. Substituting into (7), the right-hand side becomes $\tilde{O}\big(\varepsilon^{\alpha}\big)\;+\;\tilde{O}\Big(\varepsilon^{-\frac{d(d+2\alpha)}{2d+3}}\,n^{-\frac{d+3}{4d+6}}\Big)$. Morally [^2] Since the empirical process only guarantee $\sup|g_{\mathcal{D}}({\bm{u}},t)-g_{\mathcal{P}}({\bm{u}},t)|\leq\tilde{O}(n^{-1/2})$ with $n=|\mathcal{D}|$, we cannot let $g_{\mathcal{P}}(1-\varepsilon^{\star})$ less than $\tilde{O}(n^{-1/2})$, so there is a two-case discussion in Theorem 3.4., optimizing over $\varepsilon$ yields the upper bounds in Theorem 3.4, see Appendix G for details.
In the proof of the upper bound using (7), we sacrifice the shallow region $\mathbb{A}_{\varepsilon}^{d}$, but this is not just a proof artifact. Our lower bound constructions place localized ReLU atoms on disjoint spherical caps in the shallow shell where $g_{\mathcal{P}}$ is very small, so these neurons can develop large path norm at small cost. Combinations of these boundary-supported ReLUs form a family of hard-to-distinguish networks (see Construction H.4), yielding Theorem 3.5.
@@
Let $\mathcal{P}$ be any joint distribution of $({\bm{x}},y)$ where the marginal distribution of $\bm{x}$} is $\mathcal{P}_{X}(\alpha)$ and $y$ is supported on [1,1][-1,1]. Let $\mathcal{D}_{n}=\{({\bm{x}}_{j},y_{j})\}_{j=1}^{n}$ be a dataset of $n$ i.i.d. samples from $\mathcal{P}$. Let $\widehat{R}_{\mathcal{D}_{n}}(f)$ be any empirical risk estimator for the true risk $R_{\mathcal{P}}(f):=\mathbb{E}_{({\bm{x}},y)\sim\mathcal{P}}[(f({\bm{x}})-y)^{2}]$. Then,
@@
The smaller $\alpha$ is, the more mass concentrates in $\mathbb{A}_{\varepsilon}^{d}$, the more disjoint spherical caps that each carry comparable probability mass (see Figure 1(b)&(c)), the more memorization capacity the network sustain. In the extreme case where $\alpha\rightarrow 0$ and the law of $\bm{X}$} approaches $\mathrm{Uniform}(\mathbb{S}^{d-1})$, each datapoint can be isolated by a network and fitted perfectly within the BEoS regime, see Appendix I.
@@
Assume that $\{({\bm{x}}_{i},y_{i})\}_{i=1}^{n}$ is a dataset with ${\bm{x}}_{i}\in\mathbb{S}^{d-1}$ and pairwise distinct inputs. Then there exists a width $K\leq n$ network of the form (1) that interpolates the dataset and whose Hessian operator norm satisfies
@@
If we remove the hidden bias term $b$ and assume a uniform spherical distribution, the model gets harder to shatter the data, and thus yields a generalization bound of approximately $\tilde{O}(n^{-1/4})$, which is compatible with the rates in (Wu and Su, 2023).
3.2 A Unified Framework: Data Shatterability Principle
The trade-off in (6) and (7) shows that generalization is governed by the probability mass located in the $T$-deep region. Intuitively, the more mass lies in deeper regions, the stronger the implicit regularization. To quantify this geometric distribution, we introduce the following scalar summary.
Definition 3.8.
Given a distribution $\mathcal{P}_{X}$ on $\mathbb{R}^{d}$, its half-space-depth quantile function is $\Psi_{\mathcal{P}_{X}}(T):=\mathbb{P}_{{\bm{X}}\sim\mathcal{P}_{X}}(0pt({\bm{X}},\mathcal{P}_{X})\geq T)$, for $T\in[0,1/2]$. We define the half-space-depth concentration index $\mathsf{S}_{\mathrm{DQ}}(\mathcal{P}_{X}):=\big(\int_{0}^{1/2}\Psi_{\mathcal{P}_{X}}(T)\,\,\mathrm{d}T\big)^{-1}$, the reciprocal of the area under $\Psi_{\mathcal{P}_{X}}$.
(a) $\Psi_{\mathcal{P}_{X}}$ for $\mathcal{P}_{X}(\alpha)$ in $d=5$ (b) More caps ($\alpha=0.5$) (c) Less caps ($\alpha=2$)
Figure 1: Visualization of isotropic data shatterability. (a) For isotropic $\mathrm{Beta}(\alpha)$-radial distributions, larger $\alpha$ implies mass is more concentrated centrally (larger area, smaller index). (b & c) Red dots depict data points and dashed lines represent neuron boundaries. We visualize the partitioning feasibility by fixing the probability mass within each cap and comparing the number of disjoint caps that can be packed. A distribution with mass near the boundary (small $\alpha$) populates significantly more such disjoint regions than one concentrated at the center ($\alpha=2$).
We propose using this concentration index as a proxy for data shatterability, the feasibility of the data being partitioned into disjoint regions. In the isotropic case, the spherical symmetry ensures that one can always arrange a large number of disjoint activation regions (spherical caps), a quantity governed by the packing number of the high-dimensional sphere (see Lemma H.3). Consequently, the feasibility of shattering is strictly determined by the radial concentration: does each such potential cap contain sufficient probability mass? As illustrated in Figure 1(b)&(c), if we require every cap to contain a fixed amount of data, the total number of disjoint caps we can pack becomes a measure of shatterability. A large index (small $\alpha$) implies mass accumulates at the boundary, populating these pre-existing packing slots and facilitating the construction of shattering functions for our lower bounds, see Construction H.4. Conversely, a small index (large $\alpha$) starves these regions of data, limiting the noise-fitting feasibility.
A further consistency check comes from the spherical limit: for distributions on a sphere, the depth is zero everywhere. The quantile function collapses and the concentration index diverges, correctly matching the flat interpolation behavior in Theorem 3.6.
Remark 3.9 (Rectangle vs. Area).
This picture also clarifies why our upper bounds are not tight. On the $T$-deep region, we bound the entire contribution of the network by the worst-case variation over $\Omega_{T}(\mathcal{P}_{X})$, and on the shallow region we use a crude worst-case $L^{\infty}$ bound times $\mathbb{P}_{{\bm{X}}}({\bm{X}}\notin\Omega_{T}(\mathcal{P}_{X}))$, ignoring any additional regularization. Our analysis only exploits the area of a single inscribed rectangle under $T\mapsto\Psi_{\mathcal{P}_{X}}(T)$, and choosing the optimal $T^{\star}$ in (6) is equivalent to picking the largest rectangle under the curve, which discards the rest of the underlying area. This gap is exactly where potential improvements to the bounds would have to come from.
3.3 Provable Adaptation to Intrinsic Low-Dimensionality
(a) ReLUs shatter lines by knots (b) A union of $20$ lines in $\mathbb{R}^{d}$ (c) $\mathrm{Uniform}(\mathbb{B}_{1}^{d})$
Figure 2: Visualization of anisotropic data shatterability. (a) When data lies on a mixture of lines (black lines), the complex activation boundaries of ReLUs (translucent planes) reduce to a finite set of knots (colored dots). This rigidity makes the data harder to shatter than the ambient depth suggests. (b) For a union of 2020 embedded lines, the curves remain nearly identical across ambient dimensions, matching the prediction by Theorem 3.10. (c) As a comparion, the curves for $\mathrm{Uniform}(\mathbb{B}_{1}^{d})$ degenerates and reveal the curse of dimensionality predicted in (Liang et al., 2025).
We now return to anisotropic data. Here, the connection between the concentration index and generalization becomes heuristic because the geometric structure restricts the available partitioning directions. Unlike the isotropic case, where we can leverage the packing number of the ambient sphere to form disjoint regions, low-dimensional structures constrain the effective directions. For instance, on a 1D line, the high-dimensional spherical caps effectively degenerate into simple intervals or knots (Figure 2a). This collapse drastically reduces the number of possible disjoint partitions, even if the data has low ambient depth (high concentration index). Consequently, for structured data, the index serves as a conservative estimate of difficulty. Nevertheless, the principle of depth adaptation still holds: the implicit regularization adapts to the intrinsic structure rather than the ambient dimension.
Formalizing this, suppose the feature marginal is a finite mixture $\mathcal{P}_{X}=\sum_{j=1}^{J}\pi_{j}\mathcal{P}_{X,j}$. Once the depth-quantile curve $\Psi_{\mathcal{P}_{X,j}}(T)$ of each component is understood, the law of total probability yields
$\Psi_{\mathcal{P}_{X}}(T)=\mathbb{P}_{{\bm{X}}}\big(0pt({\bm{X}},\mathcal{P}_{X})\geq T\big)=\sum_{j=1}^{J}\pi_{j}\,\mathbb{P}_{{\bm{X}}}\big(0pt({\bm{X}},\mathcal{P}_{X})\geq T\mid Z=j\big).$ (11)
For any fixed $j$, the mixture cannot reduce all halfspace probabilities contributed by component $j$ by more than a factor $\pi_{j}$ (i.e., $0pt({\bm{x}},\mathcal{P}_{X})\;\geq\;\ \pi_{j}\,0pt({\bm{x}},\mathcal{P}_{X,j})$). Hence, at the level of depth-quantile curves, this shows that $\Psi_{\mathcal{P}_{X}}$ is controlled by the collection $\{\Psi_{\mathcal{P}_{X,j}}\}_{j=1}^{J}$, and the area under $\Psi_{\mathcal{P}_{X}}$ behaves like a mixture-weighted average of the areas under the $\Psi_{\mathcal{P}_{X,j}}$.
In this heuristic picture, if each component is essentially low-dimensional, then the concentration index of $\Psi_{\mathcal{P}_{X}}$ is governed by the embedded low-dimensional components and hence by the intrinsic dimension rather than the ambient one. Formally, we have the following theorem.
Theorem 3.10 (Generalization bound for mixture models).
Given a data distribution $\mathcal{P}$ on $\mathbb{R}^{d}\times\mathbb{R}$ whose feature marginal satisfies $\mathcal{P}_{X}=\sum_{j=1}^{J}\pi_{j}\mathcal{P}_{X,j}$, where each $\mathcal{P}_{X,j}$ is the uniform distribution on the unit ball in an $m$-dimensional affine subspace $V_{j}\subset\mathbb{R}^{d}$, let $\mathcal{D}=\{({\bm{x}}_{i},y_{i})\}_{i=1}^{n}$ be a dataset of $n$ i.i.d. samples drawn from $\mathcal{P}$. Then, with probability at least $1-\delta$,
$\sup_{{{\bm{\theta}}}\in{\bm{\Theta}}_{\mathrm{BEoS}}(\eta,\mathcal{D})}\mathrm{Gap}_{\mathcal{P}}(f_{{\bm{\theta}}};\mathcal{D})\;\lessapprox_{d}\;\Big(\frac{1}{\eta}-\frac{1}{2}+4M\Big)^{\frac{m}{m^{2}+4m+3}}\,M^{2}\,J^{\frac{4}{m}}\,n^{-\frac{1}{2m+4}}+M^{2}\,J\,\sqrt{\frac{1}{2n}},$ (12)
where $M\coloneqq\max\{D,\|f_{\bm{\theta}}|_{\mathbb{B}_{1}^{V}}\|_{L^{\infty}},1\}$ and $\lessapprox_{d}$ hides constants (which may depend on $d$) and logarithmic factors in $J/\delta$ and $n$.
Sketch proof of Theorem 3.10. The detailed proof is in Appendix F. The argument instantiates the above shatterability picture at the level of the BEoS-induced regularization. For each component $j$, we define a local weight function $g_{j}$ using only samples with ${\bm{X}}\in V_{j}$, and Lemma F.4 formalizes this as a uniform domination $g\gtrsim g_{j}$. Consequently, the restriction of $f_{{\bm{\theta}}}$ to $V_{j}$ has controlled $g_{j}$-weighted variation with the same order of bound. Crucially, if we restrict the network to a single $m$-dimensional subspace $V_{j}$, a neuron’s activation is governed not by its full weight vector ${\bm{w}}_{k}$, but solely by $\operatorname{proj}_{V_{j}}{\bm{w}}_{k}$, since the component of ${\bm{w}}_{k}$ that is orthogonal to $V_{j}$ is “invisible” to the data on $V_{j}$. This projection mechanism mathematically formalizes how linearly low-dimensional structures limit the model’s shatterability (as visualized by the knots in Figure 2), and thus refine the characterization of network’s effective capacity, see Theorem F.3. Moreover, some insight of this low-dimensional adaptation from the perspective of gradient dynamics is stated in Appendix B.2.
@@ 4.1 Empirical Verification of the Generalization Upper Bounds
(a) Radial concentration parameter $\alpha$ (b) Adaptation to intrinsic dimension
Figure 3: How data geometry controls generalization. (a) Fixed ambient dimension $d=5$ with isotropic Beta-radial distributions (Definition 3.2) for $\alpha\in\{0.1,0.3,1.5,5.0\}$. Larger $\alpha$ yields steeper slopes in the log–log error curve, consistent with improved rates as probability mass concentrates away from the boundary. (b) Union of $J=20$ lines ($m=1$) embedded in $\mathbb{R}^{d}$ with $d\in\{10,50,100,500\}$. The regression slopes remain nearly constant across different $d$, showing that generalization adapts to intrinsic rather than ambient dimension.
@@ 4.2 How Data Geometry Affects Representation Learning
We study how data geometry shapes the representation selected by GD at the BEoS regime through data activation rate of neurons. Given a neuron $v_{k}\phi({\bm{w}}_{k}^{\mathsf{T}}{\bm{x}}-b_{k})$ in the neural network, its data activation rate is defined as $\frac{1}{n}\sum_{i=1}^{n}\mathds{1}\{{\bm{w}}_{k}^{\top}{\bm{x}}_{i}>b_{k}\,\}$, which is exactly the probability term in the definition of the weight function $g$ in (28). Low data activation rate means the neuron fires on a small portion of the data.
@@
To verify it empirically, we compare two input distribution in $\mathbb{R}^{50}$: (i) uniform distribution on a sphere and (ii) a union of 20 lines by training ReLU networks with the same recipe and initialization . As a result, the ReLU network trained on the sphere interpolates the noisy label quickly with final true MSE 1.0249 $\approx$ noise level ($\sigma^{2}=1$), while the the ReLU network trained on a union of lines resist to overfitting with final true MSE $0.07\approx 0$ (more details appear in Appendix C). Notably, the trained representations are presented in Figure 4. In particular, GD empirically finds our lower bound construction below the edge of stability.
Figure 5: Data geometry and memorization on MNIST. Left panel: Comparison of training curves under the same ground-truth predictor with Gaussian inputs versus MNIST inputs ($n=30000$). GD on the Gaussian data set quickly interpolates, while MNIST resists overfitting for tens of thousands of steps. Right panel: Prediction error against half-space depth for MNIST samples. Shallow points (low depth) exhibit larger errors. This region refers to “highly shatterable region”.